Bonding Benchmarking

Thomas Stewart <tstewart@nibsc.ac.uk> February 28, 2006

1 Overview

I set out to experement with the bonding in linux.

2 Kit

• 2 x P4 3.0Ghz connected via gigabit capable cisco switch

• 2 x 3c900’s nics, set to 10BaseT in one box

• 1 x Broadcom gigabit nic in other box

3 Setup and Method

Straight ”modprobe bonding”, ie no link failure detection. Ran twice swapping to measure in and out. On one host: sudo dd if=/dev/zero bs=1M count=512 — sudo pv -s 512M — sudo nc host 9999 and on the other host: sudo nc -l -p 9999 — sudo pv -S 512M ¿ /dev/null

4 Results 5 Analysis

Average throuput for single 10M nic 858kB/s, range 6kB/s Theoretical max 1220kB/s, ie 30pc less

Interface In Out
eth0  859 856
eth1  862 856
bond0 1382 1607

Table 1: Cross conector for Chalsig

Average throuput for 2 bonded 10M nics 1495kB/s, range 225kB/s Theoretical max 2440kB/s, ie 38pc less